# Hydra: Breaking the Global Ordering Barrier in Multi-BFT Consensus

### v1

@@ I Introduction
Byzantine Fault Tolerant (BFT) consensus is a cornerstone of modern decentralized systems, powering diverse applications such as blockchains [9543565, QuestMarko, lens, gupta2019proof], decentralized finance (DeFi) [wang2022bft], and decentralized storage [zhu2020blockchain]. Most classical BFT protocols, including PBFT [pbft1999] and Zyzzyva [kotla2007zyzzyva], adopt a leader-based architecture, where a designated replica (also called the leader) proposes transactions and coordinates with others to reach agreement. However, as system scale grows, this traditional leader-based design becomes a key performance bottleneck [gai2021dissecting, stathakopoulou2019mir, avarikioti2020fnf, stathakopoulou2022state, gupta2021rcc, amiri2024bedrock]: the leader’s coordination workload increases linearly with the number of replicas, making it the dominant factor limiting throughput and latency.
To address this scalability bottleneck, Multi-BFT consensus has emerged as a promising direction [stathakopoulou2019mir, avarikioti2020fnf, stathakopoulou2022state, gupta2021rcc, MIR-BFT]. Multi-BFT runs multiple leader-based consensus instances (e.g., PBFT) in parallel, as illustrated in Fig. 1. Client transactions are partitioned into disjoint buckets, each handled by a separate instance, enabling concurrent agreement among multiple leaders. Each instance commits its own blocks locally, which are then merged through a global ordering phase to form a single, totally ordered ledger. This design significantly increases throughput by better utilizing available bandwidth and computational capacity.
@@
Despite its benefits, the global ordering layer–used to serialize blocks across instances–remains a major scalability barrier. While intuitively necessary for consistency, it introduces significant latency, particularly in the presence of slow instances. In practice, a single slow or faulty instance can stall the merging process, delaying confirmations for otherwise fast replicas. For example, in ISS [stathakopoulou2022state], a state-of-the-art Multi-BFT protocol, the global ordering phase alone accounts for up to 92.8% of total transaction latency with just one straggler among 16 replicas. Although recent optimizations such as Ladon [Ladon2025] and Orthrus [Orthrus2025] mitigate straggler effects, global ordering still dominates latency, contributing over 70% of total delay in Orthrus. See more details of these limitations in Sec. II.
@@
In this paper, we propose Hydra, the first Multi-BFT consensus framework that eliminates the need for global ordering. Instead of enforcing a total transaction order and executing afterward, Hydra unifies ordering and execution by leveraging transaction dependencies to expose concurrency while maintaining consistency. The key insight is that transactions operate on objects, which are independently updatable state items (e.g., accounts or contract variables). As long as operations on the same object are executed sequentially, the correctness of the overall system state is preserved. Unlike sharding—where nodes maintain only a subset of objects and must coordinate across partitions—each replica in Hydra retains a global view of all objects, enabling local coordination without expensive cross-instance protocols.
@@
We build an end-to-end prototype of Hydra in Go [golang] and conduct extensive experiments on AWS to evaluate its performance. We compare Hydra with the state-of-the-art Multi-BFT systems, including ISS [stathakopoulou2022state], RCC [gupta2021rcc], Mir-BFT [MIR-BFT], DQBFT [dqbft], Ladon [Ladon2025] and Orthrus [Orthrus2025], in terms of throughput and latency.
@@
Code availability. We will make our code publicly available once our work is published.
@@ II-B Analyzing Scalability Challenges
Table I shows experimental results to evaluate the impact of global ordering phase, which exceeds 90% of the total end-to-end latency in ISS, which is a state-of-the-art pre-determined global ordering Multi-BFT consensus. Although recent works such as Ladon [Ladon2025] and Orthrus [Orthrus2025] adopt dynamic or hybrid ordering to mitigate blocking, they still retain a global ordering barrier. Consequently, Orthrus continues to spend approximately 70% of its total latency in global ordering under the same conditions. Prior studies [stathakopoulou2022state, Orthrus2025] also obtain similar findings that under certain conditions, the global ordering stage can dominate overall latency.
@@ III-A System Model
Network model. We assume each pair of honest replicas is connected by an authenticated and reliable communication link. We adopt the partial synchrony model proposed by Dwork et al. [dwork1988consensus], commonly used in BFT consensus [pbft1999, hotstuff]. There is an established bound, denoted as $\Delta$, and an undefined Global Stabilization Time (GST). After the GST point, the delivery of any message transmitted between two honest replicas within the $\Delta$ limit is guaranteed. That is, the system behaves synchronously after the GST.
@@ III-B Data Model
Objects. We adopt an object-centric design [birrell1984implementing, open1992introduction, wollrath1996distributed, redmond1997dcom, ostrowski2008programming, weihl1988commutativity, blackshear2023sui], where objects serve as the fundamental data units processed within the system. These objects are persistent, similar to accounts, and are formally represented as a tuple: $o=(\textit{key},\textit{value})$, where key is a unique identifier, and value represents the current state of the object, which can be updated through transactions. For example, in Ethereum, each account can be treated as a distinct object, where the account address is the key and its balance is the value.
@@ IV-B Challenges and Solutions
To prevent such inconsistencies, Hydra adopts a lock-based execution model inspired by classical concurrency control mechanisms such as Two-Phase Locking (2PL) [eswaran1976notions, bernstein1987concurrency, gray1992transaction]. Each transaction must acquire locks on all of its objects before execution. Only after obtaining all locks is the transaction executed. After that, all acquired locks in the transaction are released. This ensures that no transaction is ever partially executed: either all operations complete atomically or none take effect.
Challenge 2: Conflicting ordering among instances. Different instances may impose different orders on the same set of transactions, creating deadlocks. For example, consider two transactions: $tx_{1}:A\rightarrow B$ and $tx_{2}:B\rightarrow A$. If instance 0 locks $A$ for $tx_{1}$ while instance 1 locks $B$ for $tx_{2}$, both transactions will wait indefinitely for the other’s lock. To resolve deadlocks, Hydra integrates a deadlock detection and resolution mechanism [bernstein1987concurrency]: once a transaction is confirmed but blocked on lock acquisition, the system searches for cycles in the wait-for graph and deterministically aborts one or more transactions to break the cycle.
@@ V-C Transaction Ordering
Failure detector. In Hydra, a failure detection module is integrated into the SB protocol to ensure liveness in the presence of faulty leaders. This mechanism enables replicas to replace a misbehaving or crashed leader for each SB instance, and has been widely adopted in prior BFT systems [castro2002practical, hotstuff, FireLedger, gueta2019sbft]. For example, in PBFT, when replicas suspect the leader of Byzantine behavior, they initiate a view change to elect a new one. Similarly, in Hydra, when the leader $l_{i}$ of instance $\mathit{instance}_{i}$ fails at sequence number $sn$, all honest replicas detect the failure, reach agreement on the state of $\mathit{instance}_{i}$ and the new leader $l^{\prime}_{i}$, and then restart the instance from sequence number $sn$ under the new leader.
@@ V-D Transaction Execution
Hydra ’s execution model is inspired by well-established ideas in distributed databases. The requirement that a transaction acquires all necessary object locks before execution closely echoes 2PL [eswaran1976notions, gray1992transaction, bernstein1987concurrency], which ensures atomicity and isolation by preventing other transactions from accessing locked resources. Similarly, the deadlock resolution mechanism parallels classical deadlock detection techniques that rely on identifying cycles in wait-for graphs [bernstein1987concurrency, kaveh2001deadlock].
@@ V-E Checkpoint and Garbage Collection
VI Correctness Analysis
Lemma 1.
Proof.
Lemma 2.
Proof.
Case 2: Deadlock Resolution. In this case, $tx$ is aborted due to a deadlock. By Lemma 1, all honest replicas will compute the same deadlock group $D$ by invoking $\mathsf{expandDeadlockGroup}$(txtx). All honest replicas apply the same deterministic victim selection rule (e.g., selecting the lowest-indexed transactions) in $\mathsf{selectVictims}$($D$). Since both $D$ and the selection rule are consistent across replicas, the resulting victim set is identical at all honest nodes. Therefore, if $tx\in\mathsf{selectVictims}(D)$ at one honest replica, then all replicas will also abort $tx$.
Theorem 1 (Safety).
Proof.
Lemma 3.
Proof.
Lemma 4.
If a transaction $tx$ is confirmed, it will eventually be executed.
Proof.
Theorem 2 (Liveness).
Proof.
@@ VII Performance Evaluation
In this section, we evaluate the performance of Hydra and compare it against other Multi-BFT protocols: Ladon [Ladon2025], Orthrus [Orthrus2025], ISS [stathakopoulou2022state], RCC [gupta2021rcc], Mir [MIR-BFT], and DQBFT [dqbft]. We implemented Hydra in Go [golang], using the PBFT consensus protocol [castro2002practical] as SB instances. Our evaluation aims to answer the following research questions:
@@ VII-B Performance Comparison
Fig. 3and Fig. 4 compare the throughput and latency of Hydra and other Multi-BFT protocols with one straggler and without stragglers. We evaluate the throughput and latency with 8–128 replicas in both LAN and WAN environments. We only report results with a single straggler, since performance is primarily limited by the slowest replica and adding more stragglers does not significantly change throughput or latency [Ladon2025]. For clarity, the compared protocols are categorized into three groups: pre-determined ordering schemes, including ISS [stathakopoulou2022state], Mir-BFT [MIR-BFT] and RCC [gupta2021rcc], which enforce a fixed global transaction order; dynamic ordering schemes, including DQBFT [dqbft] and Ladon [Ladon2025], which adapt ordering based on instance progress to mitigate straggler effects; and hybrid ordering scheme represented by Orthrus [Orthrus2025], which partially relaxes global ordering by leveraging fast-path confirmation for certain transactions. Following prior studies [zhang2023txallo], we set the ratio of cross-instance transactions to 12%.
Performance in WAN. Fig. 3 compares Hydra and other Multi-BFT protocols in WAN, with and without stragglers. With a straggler present, Hydra significantly outperforms pre-determined ordering schemes: with 128 replicas, throughput improves by approximately 9.0×. This is because in pre-determined ordering schemes, a slow instance delays the entire system, amplifying straggler impact. Dynamic (DQBFT, Ladon) and hybrid (Orthrus) ordering schemes alleviate this coupling but still require a global ordering phase. Consequently, Hydra continues to achieve superior performance. With 8 replicas, it achieves 47% and 38% higher throughput than Ladon and Orthrus, respectively, and with 128 replicas, Hydra surpasses DQBFT by 63%.
@@ VII-D Performance under Faults
We evaluate the performance of Hydra under crash faults in a WAN setting with 16 replicas, considering three scenarios with 0, 1, and 5 faulty replicas. Faults are injected at approximately 15s, and detected within a few seconds, triggering a view change after which performance stabilizes.
@@ VIII Discussions
Comparison with sharding protocols. Sharding-based protocols increase parallelism by partitioning both replicas and transactions across multiple shards. However, this design introduces additional coordination overhead for cross-shard transactions, as operations spanning multiple shards require multi-shard atomicity protocols. Moreover, sharding can weaken overall system security: if a shard contains a disproportionately high fraction of Byzantine replicas, the guarantees of consensus within that shard may be compromised. In contrast, Hydra maintains a global view of all transactions at each replica, allowing local coordination for both intra- and cross-instance transactions. This eliminates the need for costly cross-shard synchronization while preserving strong consistency and deterministic execution.
Distributed transaction. The key of distributed transaction processing is to ensure atomicity across partitioned data, typically through coordination protocols such as 2PC [al2017chainspace, omniledger]. While effective, these protocols incurs at least two rounds of communication and may block if the coordinator fails or if some shards are slow to respond. In contrast, Hydra operates in a fully replicated model where every replica maintains the complete object state. Cross-instance transactions are executed atomically through local locking, avoiding explicit inter-instance coordination and significantly reducing communication overhead while preserving deterministic consistency.

### v2

@@ I Introduction
Byzantine Fault Tolerant (BFT) consensus is a cornerstone of modern decentralized systems, powering diverse applications such as blockchains [33, 55, 45, 28], decentralized finance (DeFi) [56], and decentralized storage [61]. Most classical BFT protocols, including PBFT [14] and Zyzzyva [38], adopt a leader-based architecture, where a designated replica (also called the leader) proposes transactions and coordinates with others to reach agreement. However, as system scale grows, this traditional leader-based design becomes a key performance bottleneck [22, 50, 6, 52, 29, 3]: the leader’s coordination workload increases linearly with the number of replicas, making it the dominant factor limiting throughput and latency.
To address this scalability bottleneck, Multi-BFT consensus has emerged as a promising direction [50, 6, 52, 29, 51]. Multi-BFT runs multiple leader-based consensus instances (e.g., PBFT) in parallel, as illustrated in Fig. 1. Client transactions are partitioned into disjoint buckets, each handled by a separate instance, enabling concurrent agreement among multiple leaders. Each instance commits its own blocks locally, which are then merged through a global ordering phase to form a single, totally ordered ledger. This design significantly increases throughput by better utilizing available bandwidth and computational capacity.
@@
Despite its benefits, the global ordering layer–used to serialize blocks across instances–remains a major scalability barrier. While intuitively necessary for consistency, it introduces significant latency, particularly in the presence of slow instances. In practice, a single slow or faulty instance can stall the merging process, delaying confirmations for otherwise fast replicas. For example, in ISS [52], a state-of-the-art Multi-BFT protocol, the global ordering phase alone accounts for up to 92.8% of total transaction latency with just one straggler among 16 replicas. Although recent optimizations such as Ladon [43] and Orthrus [42] mitigate straggler effects, global ordering still dominates latency, contributing over 70% of total delay in Orthrus. See more details of these limitations in Sec. II.
@@
In this paper, we propose Hydra, the first Multi-BFT consensus framework that eliminates the need for global ordering. Instead of enforcing a total transaction order and executing afterward, Hydra unifies ordering and execution by leveraging transaction dependencies to expose concurrency while maintaining consistency. The key insight is that transactions operate on objects, which are independently updatable state items (e.g., accounts or contract variables). As long as operations on the same object are executed sequentially, the correctness of the overall system state is preserved. Unlike sharding—where replicas maintain only a subset of objects and must coordinate across partitions—each replica in Hydra retains a global view of all objects, enabling local coordination without expensive cross-instance protocols.
@@
We build an end-to-end prototype of Hydra in Go [25] and conduct extensive experiments on AWS to evaluate its performance. We compare Hydra with the state-of-the-art Multi-BFT systems, including ISS [52], RCC [29], Mir-BFT [51], DQBFT [4], Ladon [43] and Orthrus [42], in terms of throughput and latency.
@@ II-B Analyzing Scalability Challenges
Table I shows experimental results to evaluate the impact of global ordering phase, which exceeds 90% of the total end-to-end latency in ISS, which is a state-of-the-art pre-determined global ordering Multi-BFT consensus. Although recent works such as Ladon [43] and Orthrus [42] adopt dynamic or hybrid ordering to mitigate blocking, they still retain a global ordering barrier. Consequently, Orthrus continues to spend approximately 70% of its total latency in global ordering under the same conditions. Prior studies [52, 42] also obtain similar findings that under certain conditions, the global ordering stage can dominate overall latency.
@@ III-A System Model
Network model. We assume each pair of honest replicas is connected by an authenticated and reliable communication link. We adopt the partial synchrony model proposed by Dwork et al. [19], commonly used in BFT consensus [14, 59]. There is an established bound, denoted as $\Delta$, and an undefined Global Stabilization Time (GST). After the GST point, the delivery of any message transmitted between two honest replicas within the $\Delta$ limit is guaranteed. That is, the system behaves synchronously after the GST.
@@ III-B Data Model
Objects. We adopt an object-centric design [10, 46, 58, 48, 47, 57, 11], where objects serve as the fundamental data units processed within the system. These objects are persistent, similar to accounts, and are formally represented as a tuple: $o=(\textit{key},\textit{value})$, where key is a unique identifier, and value represents the current state of the object, which can be updated through transactions. For example, in Ethereum, each account can be treated as a distinct object, where the account address is the key and its balance is the value.
@@ IV-B Challenges and Solutions
To prevent such inconsistencies, Hydra adopts a lock-based execution model inspired by classical concurrency control mechanisms such as Two-Phase Locking (2PL) [21, 9, 26]. Each transaction must acquire locks on all of its objects before execution. Only after obtaining all locks is the transaction executed. After that, all acquired locks in the transaction are released. This ensures that no transaction is ever partially executed: either all operations complete atomically or none take effect.
Challenge 2: Conflicting ordering among instances. Different instances may impose different orders on the same set of transactions, creating deadlocks. For example, consider two transactions: $tx_{1}:A\rightarrow B$ and $tx_{2}:B\rightarrow A$. If instance 0 locks $A$ for $tx_{1}$ while instance 1 locks $B$ for $tx_{2}$, both transactions will wait indefinitely for the other’s lock. To resolve deadlocks, Hydra integrates a deadlock detection and resolution mechanism [9]: once a transaction is confirmed but blocked on lock acquisition, the system searches for cycles in the wait-for graph and deterministically aborts one or more transactions to break the cycle.
@@ V-C Transaction Ordering
Failure detector. In Hydra, a failure detection module is integrated into the SB protocol to ensure liveness in the presence of faulty leaders. This mechanism enables replicas to replace a misbehaving or crashed leader for each SB instance, and has been widely adopted in prior BFT systems [15, 59, 13, 27]. For example, in PBFT, when replicas suspect the leader of Byzantine behavior, they initiate a view change to elect a new one. Similarly, in Hydra, when the leader $l_{i}$ of instance $\mathit{instance}_{i}$ fails at sequence number $sn$, all honest replicas detect the failure, reach agreement on the state of $\mathit{instance}_{i}$ and the new leader $l^{\prime}_{i}$, and then restart the instance from sequence number $sn$ under the new leader.
@@ V-D Transaction Execution
Hydra ’s execution model is inspired by well-established ideas in distributed databases. The requirement that a transaction acquires all necessary object locks before execution closely echoes 2PL [21, 26, 9], which ensures atomicity and isolation by preventing other transactions from accessing locked resources. Similarly, the deadlock resolution mechanism parallels classical deadlock detection techniques that rely on identifying cycles in wait-for graphs [9, 36].
@@ V-E Checkpoint and Garbage Collection
VI CORRECTNESS AND ROBUSTNESS ANALYSIS
VI-A Safety and Liveness Analysis
We provide the proof sketch for safety and liveness (defined in Sec. III-D) here, while deferring the full proofs to Appendix A [44] due to space constraints.
Safety. Safety in Hydra follows from two key invariants. First, at the same delivery frontier, all honest replicas observe identical delivered transactions from each instance. This is guaranteed by the Agreement property of the underlying SB protocol. Second, deadlock detection and abort decisions are deterministic functions of these delivered transactions. Given the same set of delivered instance logs, all honest replicas compute the same deadlock group by recursively expanding conflicts. Abort decisions are then derived from this deadlock group using a deterministic victim selection rule. As a result, all honest replicas abort the same set of transactions and commit the same remaining ones.
Liveness. Liveness requires that every transaction submitted by a correct client eventually receives a response. In Hydra, this is ensured by the Termination property of the SB protocol and the epoch-based execution model. Specifically, any transaction proposed by an honest leader is eventually delivered by the SB protocol. Transactions that cannot be committed due to conflicts or deadlocks are deterministically aborted and reinserted for execution in subsequent epochs. The epoch mechanism bounds lock holding time and prevents transactions from being blocked indefinitely.
VI-B Impact of Byzantine Behaviors
We analyze how Hydra behaves under Byzantine behaviors that deliberately target its locking mechanisms and cross-instance coordination.
Lock manipulation by Byzantine leaders. A Byzantine leader could manipulate locking behavior from two aspects.
(i) Delaying locked trasactions. A Byzantine leader may intentionally delay proposing a cross-instance transaction in one instance after it has already been proposed and has acquired locks in another instance. For example, for a transaction spanning instances $ins_{1}$ and $ins_{2}$, a Byzantine leader in $ins_{2}$ may withhold or delay its proposal, causing locks acquired in $ins_{1}$ to be held for an extended period. This behavior increases blocking for other transactions accessing the same objects and degrades throughput.
(ii) Creating deadlocks. A Byzantine leader may also attempt to induce deadlocks by proposing conflicting orderings across instances. For instance, two cross-instance transactions $tx_{1}$ and $tx_{2}$ may be ordered as $tx_{1}\rightarrow tx_{2}$ in $ins_{1}$, while a Byzantine leader in $ins_{2}$ intentionally proposes them in the reverse order $tx_{2}\rightarrow tx_{1}$, creating a deadlock.
Adversarial transaction injection by malicious clients. Malicious clients may deliberately generate multiple transactions involving the same set of objects and submit them to different instances in adversarial orders, intentionally creating deadlocks across instances and amplifying the likelihood of cross-instance deadlocks.
Impact analysis of Byzantine behaviors. The above adversarial behaviors target the liveness of the system rather than safety. Byzantine leaders or malicious clients may delay locked transactions or induce cross-instance deadlocks, thereby degrading throughput and increasing latency. However, the impact of such adversarial behaviors is bounded. First, deadlock resolution ensures that any induced deadlock is consistently identified system-wide. Once detected, the deadlock resolution protocol deterministically aborts a bounded set of transactions, preventing indefinite blocking. Second, the epoch-based execution model enforces an upper bound on lock holding time: transactions that fail to make progress are aborted and retried in subsequent epochs, eliminating the possibility of persistent lock monopolization by Byzantine behavior. As a result, while adversaries can temporarily reduce system efficiency, they cannot cause unbounded delays, inconsistent execution, or divergence among honest replicas.
@@ VII Performance Evaluation
In this section, we evaluate the performance of Hydra and compare it against other Multi-BFT protocols: Ladon [43], Orthrus [42], ISS [52], RCC [29], Mir [51], and DQBFT [4]. We implemented Hydra in Go [25], using the PBFT consensus protocol [15] as SB instances. Our evaluation aims to answer the following research questions:
@@
  • $\bullet$
    Q4: How does Hydra ’s memory usage scale with the number of instances? (Sec. VII-E)
@@ VII-B Performance Comparison
Fig. 3and Fig. 4 compare the throughput and latency of Hydra and other Multi-BFT protocols with one straggler and without stragglers. We evaluate the throughput and latency with 8–128 replicas in both LAN and WAN environments. We only report results with a single straggler, since performance is primarily limited by the slowest replica and adding more stragglers does not significantly change throughput or latency [43]. For clarity, the compared protocols are categorized into three groups: pre-determined ordering schemes, including ISS [52], Mir-BFT [51] and RCC [29], which enforce a fixed global transaction order; dynamic ordering schemes, including DQBFT [4] and Ladon [43], which adapt ordering based on instance progress to mitigate straggler effects; and hybrid ordering scheme represented by Orthrus [42], which partially relaxes global ordering by leveraging fast-path confirmation for certain transactions. Following prior studies [60], we set the ratio of cross-instance transactions to 12%.
Performance in WAN. Fig. 3 compares Hydra and other Multi-BFT protocols in WAN, with and without stragglers. With a straggler present, Hydra significantly outperforms pre-determined ordering schemes: with 128 replicas, throughput improves by approximately 9.0×. This is because in pre-determined ordering schemes, a slow instance delays the entire system, amplifying straggler impact. Dynamic (DQBFT, Ladon) and hybrid (Orthrus) ordering schemes alleviate this coupling but still require a global ordering phase. For DQBFT, throughput decreases with more replicas because its single ordering instance becomes a bottleneck as the system scales. Consequently, Hydra continues to achieve superior performance. With 8 replicas, it achieves 47% and 38% higher throughput than Ladon and Orthrus, respectively, and with 128 replicas, Hydra surpasses DQBFT by 63%.
@@ VII-D Performance under Faults
Crash faults. We evaluate the performance of Hydra under crash faults in a WAN setting with 16 replicas, considering three scenarios with 0, 1, and 5 faulty replicas. Faults are injected at approximately 15s, and detected within a few seconds, triggering a view change after which performance stabilizes.
@@
Table II: Throughput and latency under different adversarial scenarios. BL: Byzantine Leader, MC: Malicious Clients.
Metric / Scenario Failure-free MC BL BL + MC
Throughput (ktps) 61.50 57.01 60.13 52.86
Latency (s) 2.69 3.26 3.02 3.60
Byzantine faults. Table II summarizes system performance under different adversarial scenarios, including high-contention workloads generated by malicious clients (MC), Byzantine leader behavior (BL), and the combination of both (BL+MC). In the BL scenario, we inject a single Byzantine leader among 16 replicas, which deliberately prolongs lock holding and proposes adversarial transaction orderings to induce cross-instance deadlocks. In the MC scenario, malicious clients generate high-contention workloads by repeatedly submitting transactions involving the same set of 20 objects to different instances in adversarial orders, intentionally inducing cross-instance deadlocks and increasing coordination overhead. Each scenario captures realistic adversarial strategies that may increase lock contention or coordination delays, as discussed in Section VI-B.
As the results show, although adversarial behaviors can reduce throughput and increase latency, the impact on Hydra is limited. Malicious clients submitting high-contention transactions reduce throughput by 7.3% and increase latency by 21.2%. Byzantine leaders attempting to manipulate locks or propose conflicting orders cause a smaller throughput reduction of 2.2% and a latency increase of 12.3%. When both adversarial behaviors are combined (BL+MC), throughput decreases by 14.0% and latency increases by 33.8%. These results indicate that Hydra’s locking and deadlock handling mechanisms effectively contain the effects of adversarial manipulation, restricting the impact to performance degradation.
VII-E Memory Usage
Table III: Memory usage and utilization under different numbers of instances. Each instance: 8 vCPUs and 16 GB RAM.
# Instances 8 16 32 64
Memory Usage (MB) 11046 10648 10549 10560
Utilization 67.4% 65.0% 64.4% 64.5%
[…]
This behavior is attributed to Hydra’s epoch-based processing. At the end of each epoch, replicas perform checkpointing, truncate committed logs, and apply garbage collection to reclaim memory. As a result, the in-memory state maintained by each replica is bounded by the most recent checkpoint and does not grow unbounded over time, ensuring sustainable memory usage for long-running deployments.
@@ VIII Discussions
Comparison with sharding protocols. Sharding-based protocols increase parallelism by partitioning both replicas and transactions across multiple shards. However, this design introduces additional coordination overhead for cross-shard transactions, as operations spanning multiple shards require multi-shard atomicity protocols. Hydra ’s core advantage lies in its communication-free cross-instance execution model, which fundamentally differentiates it from existing approaches. This point is further elaborated in the Appendix C of [44].
Distributed transaction. The key of distributed transaction processing is to ensure atomicity across partitioned data, typically through coordination protocols such as 2PC [1, 37]. While effective, these protocols incurs at least two rounds of communication and may block if the coordinator fails or if some shards are slow to respond. In contrast, Hydra operates in a fully replicated model where every replica maintains the complete object state. Cross-instance transactions are executed atomically through local locking, avoiding explicit inter-instance coordination and significantly reducing communication overhead while preserving deterministic consistency.
Transaction partition strategies. Hydra does not assume a specific transaction partitioning algorithm, but instead is designed to operate correctly and efficiently under a wide range of existing partitioning strategies. For example, a simple and widely adopted approach assigns objects to instances by hashing their keys, i.e., ${insIndex}=\mathsf{Hash}({key})\bmod m$. A transaction is then routed to the instances responsible for the objects it accesses. This simple strategy achieves good load balance in expectation, but it may increase the frequency of cross-instance transactions when objects that frequently interact are mapped to different instances.
To mitigate this effect, Hydra can naturally benefit from prior work such as TxAllo [60], which reduces the frequency of cross-instance transactions while maintaining a balanced load across instances. By co-locating frequently interacting objects when possible, such approaches lower cross-instance coordination overhead and deadlock probability, thereby improving overall throughput.
Importantly, these partitioning strategies are orthogonal to the core design of Hydra. The system’s correctness guarantees hold under any deterministic partitioning scheme agreed upon by all replicas.